# DeepShade: Enable Shade Simulation by Text-conditioned Image Generation

### v1

@@ 6 Demonstration
Besides the quantitative analysis in experiments, it is important to demonstrate the real-world impact of the work. Thus, we design a proof-of-concept demo as shown in Figure. 4 using a subarea of Arizona State University. In this demo, we tackle the problem of integrating the shade ratio as a factor when making the routing suggestions. The input is a skeleton image containing the building outline, extracted from a satellite image of the interested area. Then, based on the time for planning, we describe as the text prompt $T$ together with the image that is processed to $x^{cond}$ as in Eq. 5, the shape map will be generated as output, it will be used for shade ratio calculation by overlaying the shade map with the road using longitude and latitude ranges. Given the shade ratio, the planning is made by jointly considering the user’s preference (weight) on shade and distance by a variant of the Dijkstra algorithm. The green shows a more shaded plan, while the red means the shortest path. This demo reveals the potential of a real-world application in this direction; shade-involved planning is crucial for areas that suffer from extreme heat waves, and could possibly help decrease heat stroke cases, improve the health of outdoor people.

### v2

@@ 6 Demonstration
Besides the quantitative analysis in experiments, it is important to demonstrate the real-world impact of the work. Thus, we design a proof-of-concept demo as shown in Figure. 4 using a subarea of Arizona State University. In this demo, we tackle the problem of integrating the shade ratio as a factor when making the routing suggestions. The input is a skeleton image containing the building outline, extracted from a satellite image of the interested area. Then, based on the time for planning, we describe as the text prompt $T$ together with the image that is processed to $x^{cond}$ as in Eq. 5, the shape map will be generated as output, it will be used for shade ratio calculation by overlaying the shade map with the road using longitude and latitude ranges. Given the shade ratio, the planning is made by jointly considering the user’s preference (weight) on shade and distance by a variant of the Dijkstra algorithm. The green shows a more shaded plan, while the red means the shortest path. This demo reveals the potential of a real-world application; given that shade-involved planning is crucial for areas that suffer from extreme heat waves, this demonstration shows a way that could possibly help decrease heat stroke cases and improve the health of outdoor people.